A particle is moving with the given data. Find the position of the particle. a(t) = 2t + 3, 5(0) = 5, V(0) = -2 s(t) =
Answer #1

alt) = 2t+3 vetr = Jacly at - S 2t+3 dt = +²+ 3t to vor= -2 -2a ototo I vet) = x² + 34-2 Seth = Sucts are a t² + 3t - 2) at -(0) = 5= oto-0=0 schef + 3-20+5 ]